# A Survey on LLM-as-a-Judge

### v4

@@ 7.1.2. Social Intelligence
As LLM capabilities advance, machines are increasingly taking on tasks once deemed exclusive to humans, particularly in context-specific domains. A notable area is social intelligence, where models must navigate complex social scenarios involving cultural values, ethical principles, and social impacts. For instance, Xu et al. (2024) assess the social intelligence of LLMs, noting that despite significant progress, these models still fall short compared to their academic problem-solving abilities. Similarly, Zhou et al. (2023b) introduce SOTOPIA and SOTOPIA-EVAL to simulate intricate social interactions among LLM agents and evaluate their social intelligence. In their study, GPT-4 serves as a stand-in for human judgment, assessing goal completion, financial management, and relationship preservation within these simulated interactions.
@@ 7.1.3. Multi-Modal
In the field of multi-modal AI, benchmarks have been created to assess LLM-based systems that function across text and vision modalities. These benchmarks have enabled the evaluation of tasks such as image captioning and mathematical reasoning, where LLMs aligned with human preferences in pairwise comparisons but performed poorly in scoring and batch ranking (Chen et al., 2024b). For Chinese multi-modal alignment, benchmarks have identified challenges in coherence and reasoning, leading to the proposal of a calibrated evaluation model that achieves greater consistency than existing systems (Wu et al., 2024). Furthermore, advancements in multi-modal and multi-agent systems have been reviewed, emphasizing collaboration mechanisms to improve rationality and minimize biases (Jiang et al., 2024). Xiong et al. (2024) investigate the use of LLM-as-a-judge to assess the performance of multimodal models, offering both a final score and the rationale behind the evaluations to enhance transparency and consistency. Chen et al. (2024f) introduce the first benchmark for the automatic evaluation of LVLMs, focusing on self-driving corner cases. Their findings indicate that evaluations conducted by LLMs-as-judges align more closely with human preferences compared to those performed by LVLM-as-judges.
@@ 7.2.1. Finance
In the context of LLM-as-a-judge applications within finance, expert knowledge is crucial for domain-specific evaluations. Current research can be divided into two areas: one focuses on designing LLM-based evaluators that leverage expert knowledge for specific tasks. For instance, Brief et al. (2024) conducted a case study on multi-task fine-tuning in finance to enhance LLM performance (Brief et al., 2024), while Yu et al. (2024) introduced FinCon, a multi-agent system that uses conceptual verbal reinforcement to improve financial decision-making (Yu et al., 2024a). The second area of research aims to provide benchmarks to evaluate and enhance the understanding of domain-specific knowledge by LLMs. These benchmarks include UCFE (Yang et al., 2024a) based on user feedback, IndoCareer: a dataset of professional exam questions (Koto, 2024), and AI-generated domain-specific evaluation sets (Raju et al., 2024). In quantitative investment, LLM-as-a-judge approaches have demonstrated value in refining and enhancing LLM-generated trading signals. (Wang et al., 2024g) proposes a two-layer architecture for generating self-improving trading signals. Their system employs a dual-LLM setup in the inner loop, where one LLM generates trading ideas while another serves as a judge to evaluate and refine them. The outer loop incorporates an additional LLM judge that provides comprehensive reviews based on quantitative metrics such as information coefficient and Sharpe ratio, ensuring trading signals meet rigorous performance standards.
@@ 7.2.3. AI for Science
LLMs have demonstrated notable potential in scientific fields, especially in areas like medical question-answering and mathematical reasoning, where they serve as evaluators to improve accuracy and consistency. In the medical field, studies by Brake et al. (2024) and Krolik et al. (2024) showed that models like LLaMA2 can assess clinical notes and Q&A responses with a level of accuracy approaching that of human experts (Brake and Schaaf, 2024; Krolik et al., 2024). This approach leverages prompt engineering to embed expert knowledge, enabling LLMs to handle complex, nuanced information, which provides a reliable first-line assessment that lessens the load on human experts.
In mathematical reasoning, reinforcement learning (RL) and cooperative reasoning methods further enhance LLM’s capability as an evaluator, especially for theorem-proving tasks (Lu et al., 2023). For example, WizardMath was introduced by employing RL through step-by-step feedback to refine reasoning in mathematical tasks (Luo et al., 2023b). Zhu et al. (2023) proposed a Cooperative Reasoning (CoRe) framework that combines generation and verification to mimic human-like dual-process reasoning, enhancing the model’s problem-solving accuracy (Zhu et al., 2023c). Additionally, Lu et al. (2023) developed MathVista, a benchmark for evaluating mathematical reasoning in visual contexts, which assesses LLMs like GPT-4V on tasks involving mathematical reasoning with visual components (Lu et al., 2023). These methods highlight the value of combining RL, cooperative reasoning, and prompt engineering in improving LLMs’ evaluative and reasoning skills across mathematical reasoning.
@@ 8.2. Robustness
Despite LLM’s superior power, it is found prone to adversarial attacks (Shen et al., 2023; Jiang et al., 2023; Zou et al., 2023), under which LLMs can be induced to generate harmful content. While existing works on LLM attacks mainly focus on NLG tasks, more attacks on LLM-as-a-judge are relatively under-explored (Chen et al., 2024a). This means that we will face some robustness challenges when using LLM-as-a-Judge, and these risks are unknown.
@@ 9. Future Work
In the AI era, LLM-as-a-Judge systems are increasingly demonstrating their potential to assist or even replace human judgment across a wide range of professional domains. Many roles inherently require the ability to evaluate, assess, or adjudicate complex scenarios, and LLMs, with their advanced data processing and pattern recognition capabilities, are well-suited to support or enhance these tasks. As shown in Figure 16, LLMs can serve as versatile evaluators in diverse fields. For instance, writers can leverage LLMs to assess the viability and originality of creative ideas by analyzing narrative structures and market trends; doctors can use LLMs to diagnose conditions and predict outcomes by processing medical records and imaging data; quantitative analysts can employ LLMs to forecast market movements and assess risks by identifying patterns in financial data; and judges can rely on LLMs to interpret laws and precedents, aiding in the adjudication of legal cases.
While LLMs excel in scalable and flexible evaluation, they do have limitations. Future work should focus on addressing these limitations while exploring new applications and improving the reliability, fairness, and adaptability of LLM-as-a-Judge systems to ensure their alignment with societal values and professional standards.
@@ 9.1. More Reliable LLM-as-a-Judge
As highlighted in our Formulation (§ 2) and Strategy (§ 3), LLMs are probabilistic models that require extensive research and optimization to enhance their reliability as judges. Although current methods have improved the reliability of LLM-as-a-Judge, many challenges, including adaptability and robustness, remain unresolved. To enable probabilistic models to deliver evaluations closely aligned with real-world scenarios, future research should prioritize refining and implementing LLM-as-a-Judge across the evaluation pipeline.
There is considerable potential for improving reliability in various aspects, including in-context learning, model selection, post-processing techniques, and the overall evaluation framework for LLM-as-a-Judge. These efforts should prioritize not only enhancing the reliability of assessments but also developing methodologies to systematically evaluate and validate the robustness of these assessments. Furthermore, the establishment of comprehensive evaluation benchmarks and interpretable analytical tools will be crucial for assessing and improving the reliability of LLM evaluators.
Finally, the uncertain and evolving nature of robustness risks underscores the necessity of proactive mitigation strategies. These strategies should include the development of adversarial training techniques tailored to judgment tasks, the integration of robust uncertainty quantification methods, and the implementation of human-in-the-loop systems to oversee critical decisions. By addressing these challenges, we can build more resilient and dependable systems capable of maintaining high levels of reliability even under adversarial conditions.
@@ 9.2. LLM-as-a-Judge for Data Annotation
In contrast, LLM-as-a-judge is a general technique where you use LLM to approximate human labeling. When you ask an LLM to assess qualities like ”faithfulness to source,” ”correctness,” or ”helpfulness,” you define what these terms mean in the evaluation prompt and rely on the semantic relationships the LLM learned from training data. Despite its wide applications, data annotation poses significant challenges for current machine-learning models due to the complexity, subjectivity, and diversity of data. This process requires domain expertise and is resource-intensive, particularly when manually labeling large datasets. Advanced LLMs such as GPT-4 (OpenAI, 2023a), Gemini (Google, 2023), and LLaMA-2 (Touvron et al., 2023c) offer a promising opportunity to revolutionize data annotation. LLMs serve as more than just tools but play a crucial role in improving the effectiveness and precision of data annotation. Their ability to automate annotation tasks (Zhang et al., 2023d), ensure consistency across large volumes of data, and adapt through fine-tuning or prompting for specific domains (Song et al., 2024a; Ma et al., 2024), significantly mitigates the challenges encountered with traditional annotation methods, setting a new standard for what is achievable in the realm of NLP.
@@ 9.3. MLLM-as-a-Judge
AI systems are evolving into highly versatile and multifunctional entities. Traditionally, specialized models were required for distinct language processing tasks, such as sentiment analysis, syntactic parsing, and dialogue modeling. However, large language models (LLMs) have demonstrated competence across these tasks using a single set of weights (Srivastava et al., 2022). Similarly, advancements are being made toward unified systems capable of processing multiple data modalities. Instead of employing distinct architectures for processing text, audio, and images, recent models like GPT-4o (OpenAI, 2023b), Gemini (Google, 2023), and LLaVA (Liu et al., 2023b) integrate these capabilities within a single framework. These developments highlight a growing trend toward unification in the structure and functionality of AI systems, which extends to the emerging paradigm of LLM-as-a-Judge.
Currently, MLLM-as-a-Judge frameworks (Chen et al., 2024c) are emerging for evaluating models. However, research exploring how MLLM-as-a-Judge could be applied to the evaluation of data or agents remains limited. Beyond model evaluation, MLLM-as-a-Judge, much like LLM-as-a-Judge, is envisioned to have the capability to assess or annotate data, function as a Reward Model, or serve as a Verifier within intermediate reasoning processes. These expanded roles would allow MLLM-as-a-Judge to contribute more broadly to the AI pipeline.
@@ 10. Conclusion
LLM-as-a-Judge represents an LLM-based paradigm for evaluation, offering a compelling alternative to traditional methods that rely on either human experts or quantitative assessments. By leveraging the powerful capabilities of LLMs, this framework excels in tasks such as scoring, rating, and evaluation, offering exceptional scalability, adaptability, and precision. These qualities align directly with the growing demand for efficient and robust evaluation systems across diverse fields.
However, fully realizing the potential of LLM-as-a-Judge requires addressing significant challenges, particularly those concerning reliability. The evaluation process must account for factors such as consistency, bias mitigation, and contextual adaptability, ensuring that the system operates under well-defined rules while delivering human-aligned and objective results. Each stage of development—from dataset curation and model fine-tuning to evaluation protocol standardization—demands careful consideration to uphold reliability. Simply deploying LLM-as-a-Judge does not guarantee outcomes that are both accurate and aligned with evaluation needs. Building reliable systems requires systematic efforts to evaluate and refine their outputs, while addressing gaps in definitions, practices, and research.
This survey has provided a comprehensive review of the LLM-as-a-Judge landscape, offering insights into its current state and practical strategies for improvement. By addressing these challenges with robust methodologies and innovative benchmarks, LLM-as-a-Judge can evolve into a highly reliable and practical tool for diverse applications. Ultimately, while the current capabilities of LLMs are impressive, their effectiveness as evaluators depends on thoughtful design and systematic refinement. We hope this work serves as a foundation for further research, guiding the development of more reliable LLM-as-a-Judge systems and enabling their broader deployment across academic, industrial, and societal domains.

### v5

@@ 7.1.2. Social Intelligence
As LLM capabilities advance, machines are increasingly taking on tasks once deemed exclusive to humans, particularly in context-specific domains. A notable area is social intelligence, where models must navigate complex social scenarios involving cultural values, ethical principles, and social impacts. For instance, Xu et al. (2024a) assess the social intelligence of LLMs, noting that despite significant progress, these models still fall short compared to their academic problem-solving abilities. Similarly, Zhou et al. (2023b) introduce SOTOPIA and SOTOPIA-EVAL to simulate intricate social interactions among LLM agents and evaluate their social intelligence. In their study, GPT-4 serves as a stand-in for human judgment, assessing goal completion, financial management, and relationship preservation within these simulated interactions.
@@ 7.1.3. Multi-Modal
In the field of multi-modal AI, benchmarks have been created to assess LLM-based systems that function across text and vision modalities. These benchmarks have enabled the evaluation of tasks such as image captioning and mathematical reasoning, where LLMs aligned with human preferences in pairwise comparisons but performed poorly in scoring and batch ranking (Chen et al., 2024b). For Chinese multi-modal alignment, benchmarks have identified challenges in coherence and reasoning, leading to the proposal of a calibrated evaluation model that achieves greater consistency than existing systems (Wu et al., 2024). Furthermore, advancements in multi-modal and multi-agent systems have been reviewed, emphasizing collaboration mechanisms to improve rationality and minimize biases (Jiang et al., 2024). Xiong et al. (2024) investigate the use of LLM-as-a-judge to assess the performance of multimodal models, offering both a final score and the rationale behind the evaluations to enhance transparency and consistency. Chen et al. (2024d) introduce the first benchmark for the automatic evaluation of LVLMs, focusing on self-driving corner cases. Their findings indicate that evaluations conducted by LLMs-as-judges align more closely with human preferences compared to those performed by LVLM-as-judges.
Figure 16. Scenarios where the role that can be assisted by LLM-as-a-Judge is applied.
@@ 7.2.1. Finance
In the context of LLM-as-a-judge applications within finance, expert knowledge is crucial for domain-specific evaluations. Current research can be divided into two areas: one focuses on designing LLM-based evaluators that leverage expert knowledge for specific tasks. For instance, Brief et al. (2024) conducted a case study on multi-task fine-tuning in finance to enhance LLM performance (Brief et al., 2024), while Yu et al. (2024) introduced FinCon, a multi-agent system that uses conceptual verbal reinforcement to improve financial decision-making (Yu et al., 2024a). The second area of research aims to provide benchmarks to evaluate and enhance the understanding of domain-specific knowledge by LLMs. These benchmarks include UCFE (Yang et al., 2024a) based on user feedback, IndoCareer: a dataset of professional exam questions (Koto, 2024), and AI-generated domain-specific evaluation sets (Raju et al., 2024). In quantitative investment, LLM-as-a-judge approaches have demonstrated value in refining and enhancing LLM-generated trading signals. (Wang et al., 2024g) (Figure 17) proposes a two-layer architecture for generating self-improving trading signals. Their system employs a dual-LLM setup in the inner loop, where one LLM generates trading ideas while another serves as a judge to evaluate and refine them. The outer loop incorporates an additional LLM judge that provides comprehensive reviews based on quantitative metrics such as information coefficient and Sharpe ratio, ensuring trading signals meet rigorous performance standards.
@@
Figure 17. Illustration of using dual-LLM iterative feedback loop for alpha generation in finance. Figure adapted from Wang et al. (2024g).
@@ 7.2.3. AI for Science
LLMs have demonstrated notable potential in scientific fields Zhou et al. (2024a); Tang et al. (2024b); Zhao et al. ([n. d.]), especially in areas like medical question-answering and mathematical reasoning, where they serve as evaluators to improve accuracy and consistency. In the medical field, studies by Brake et al. (2024) and Krolik et al. (2024) showed that models like LLaMA2 can assess clinical notes and Q&A responses with a level of accuracy approaching that of human experts (Brake and Schaaf, 2024; Krolik et al., 2024). This approach leverages prompt engineering to embed expert knowledge, enabling LLMs to handle complex, nuanced information, which provides a reliable first-line assessment that lessens the load on human experts.
In mathematical reasoning, reinforcement learning (RL) and cooperative reasoning methods further enhance LLM’s capability as an evaluator, especially for theorem-proving tasks (Lu et al., 2023). For example, WizardMath was introduced by employing RL through step-by-step feedback to refine reasoning in mathematical tasks (Luo et al., 2023). Zhu et al. (2023) proposed a Cooperative Reasoning (CoRe) framework that combines generation and verification to mimic human-like dual-process reasoning, enhancing the model’s problem-solving accuracy (Zhu et al., 2023b). Additionally, Lu et al. (2023) developed MathVista, a benchmark for evaluating mathematical reasoning in visual contexts, which assesses LLMs like GPT-4V on tasks involving mathematical reasoning with visual components (Lu et al., 2023). These methods highlight the value of combining RL, cooperative reasoning, and prompt engineering in improving LLMs’ evaluative and reasoning skills across mathematical reasoning.
@@ 8.2. Robustness
Despite LLM’s superior power, it is found prone to adversarial attacks (Shen et al., 2023; Jiang et al., 2023a; Zou et al., 2023), under which LLMs can be induced to generate harmful content. While existing works on LLM attacks mainly focus on NLG tasks, more attacks on LLM-as-a-judge are relatively under-explored (Chen et al., 2024a). This means that we will face some robustness challenges when using LLM-as-a-Judge, and these risks are unknown.
@@ 9. Future Work
In the AI era (Xu et al., 2024b), LLM-as-a-Judge systems are increasingly demonstrating their potential to assist or even replace human judgment across a wide range of professional domains. Many roles inherently require the ability to evaluate, assess, or adjudicate complex scenarios, and LLMs, with their advanced data processing and pattern recognition capabilities, are well-suited to support or enhance these tasks. As shown in Figure 16, LLMs can serve as versatile evaluators in diverse fields (Wang et al., 2023d). For instance, writers can leverage LLMs to assess the viability and originality of creative ideas by analyzing narrative structures and market trends; doctors can use LLMs to diagnose conditions and predict outcomes by processing medical records and imaging data (Zhang et al., 2023b; Tang et al., 2024b; Zhong et al., 2023); quantitative analysts can employ LLMs to forecast market movements and assess risks by identifying patterns in financial data; and judges can rely on LLMs to interpret laws and precedents, aiding in the adjudication of legal cases. While LLMs excel in scalable and flexible evaluation, they do have limitations. Future work should focus on addressing these limitations while exploring new applications and improving the reliability, fairness, and adaptability of LLM-as-a-Judge systems to ensure their alignment with societal values and professional standards.
@@ 9.1. More Reliable LLM-as-a-Judge
As highlighted in our Formulation (§ 2) and Strategy (§ 3), LLMs are probabilistic models that require extensive research and optimization to enhance their reliability as judges. Although current methods have improved the reliability of LLM-as-a-Judge, many challenges, including adaptability and robustness, remain unresolved. To enable probabilistic models to deliver evaluations closely aligned with real-world scenarios, future research should prioritize refining and implementing LLM-as-a-Judge across the evaluation pipeline. There is considerable potential for improving reliability in various aspects, including in-context learning, model selection, post-processing techniques, and the overall evaluation framework for LLM-as-a-Judge. These efforts should prioritize not only enhancing the reliability of assessments but also developing methodologies to systematically evaluate and validate the robustness of these assessments. Furthermore, the establishment of comprehensive evaluation benchmarks and interpretable analytical tools will be crucial for assessing and improving the reliability of LLM evaluators. Finally, the uncertain and evolving nature of robustness risks underscores the necessity of proactive mitigation strategies. These strategies should include the development of adversarial training techniques tailored to judgment tasks, the integration of robust uncertainty quantification methods, and the implementation of human-in-the-loop systems to oversee critical decisions. By addressing these challenges, we can build more resilient and dependable systems capable of maintaining high levels of reliability even under adversarial conditions.
@@ 9.2. LLM-as-a-Judge for Data Annotation
In contrast, LLM-as-a-judge is a general technique where you use LLM to approximate human labeling. When you ask an LLM to assess qualities like ”faithfulness to source,” ”correctness,” or ”helpfulness,” you define what these terms mean in the evaluation prompt and rely on the semantic relationships the LLM learned from training data. Despite its wide applications, data annotation poses significant challenges for current machine-learning models due to the complexity, subjectivity, and diversity of data. This process requires domain expertise and is resource-intensive, particularly when manually labeling large datasets. Advanced LLMs such as GPT-4 (OpenAI, 2023), Gemini (Google, 2023), and LLaMA-2 (Touvron et al., 2023c) offer a promising opportunity to revolutionize data annotation. LLMs serve as more than just tools but play a crucial role in improving the effectiveness and precision of data annotation. Their ability to automate annotation tasks (Zhang et al., 2023e), ensure consistency across large volumes of data, and adapt through fine-tuning or prompting for specific domains (Song et al., 2024a; Ma et al., 2024), significantly mitigates the challenges encountered with traditional annotation methods, setting a new standard for what is achievable in the realm of NLP.
@@ 9.3. MLLM-as-a-Judge
AI systems are evolving into highly versatile and multifunctional entities (DATA, 2024). Traditionally, specialized models were required for distinct language processing tasks, such as sentiment analysis, syntactic parsing, and dialogue modeling. However, large language models (LLMs) have demonstrated competence across these tasks using a single set of weights (Srivastava et al., 2022). Similarly, advancements are being made toward unified systems capable of processing multiple data modalities. Instead of employing distinct architectures for processing text, audio, and images, recent models like GPT-4o (OpenAI, 2023), Gemini (Google, 2023), and LLaVA (Liu et al., 2023c) integrate these capabilities within a single framework. These developments highlight a growing trend toward unification in the structure and functionality of AI systems, which extends to the emerging paradigm of LLM-as-a-Judge.
Currently, MLLM-as-a-Judge frameworks (Chen et al., 2024b) are emerging for evaluating models. However, research exploring how MLLM-as-a-Judge could be applied to the evaluation of data or agents remains limited. Beyond model evaluation, MLLM-as-a-Judge, much like LLM-as-a-Judge, is envisioned to have the capability to assess or annotate data, function as a Reward Model, or serve as a Verifier within intermediate reasoning processes. These expanded roles would allow MLLM-as-a-Judge to contribute more broadly to the AI pipeline.
@@ 10. Conclusion
LLM-as-a-Judge has emerged as a promising paradigm for automated evaluation, offering scalability and adaptability that surpass traditional expert-driven or metric-based methods. By leveraging the reasoning capabilities of large language models, this framework excels in tasks such as text quality assessment, model evaluation, and automated data annotation. It is particularly valuable for large-scale, efficient, and adaptable evaluation. Its ability to process diverse content formats and integrate domain-specific knowledge makes it particularly well-suited for applications in education, peer review, and decision-making systems.
Despite these strengths, several challenges must be addressed to fully realize its potential. Ensuring reliability remains a key issue, because probabilistic outputs can introduce inconsistencies, overconfidence, and biases inherited from training data. Although techniques RLHF have improved alignment with human judgment, they do not eliminate all sources of subjectivity. Moreover, ensuring robustness is another critical concern. LLM-as-a-Judge can be susceptible to adversarial prompt manipulation and contextual framing biases, potentially causing unintended or unreliable evaluations. Finally, generalization across domains and modalities remains a significant hurdle, as current models struggle with evaluating multi-modal inputs, reasoning over structured data, and adapting to domain-specific evaluation standards.
To address these challenges, future research should focus on three key areas. First, improving reliability requires advancements in self-consistency mechanisms, uncertainty calibration, and bias mitigation techniques, ensuring that models provide stable and well-calibrated judgments. Second, enhancing robustness involves developing adversarial-resistant evaluation frameworks and refining prompt engineering methodologies, reducing sensitivity to context variations. Third, expanding generalization capabilities calls for advancing multi-modal reasoning, integrating structured knowledge representations, and refining domain-adaptive learning strategies, allowing models to handle diverse evaluation scenarios more effectively.
Ultimately, LLM-as-a-Judge is poised to become an integral component of next-generation evaluation systems, augmenting human expertise rather than replacing it. By addressing the challenges of reliability, robustness, and generalization, we can create more trustworthy, adaptive, and comprehensive evaluators, paving the way for their adoption across scientific research, education, industry, and beyond.